# SceneDreamer360: Text-Driven 3D-Consistent Scene Generation with Panoramic Gaussian Splatting

### v1

@@ Introduction
Step 1: ControlNet-Tile Super-Resolution Reconstruction The ControlNet-Tile model divides the panorama into distinct tiles through semantic recognition. Each tile is analyzed by the model to determine which details to add and which parts to repair. Based on tiling, ControlNet-Tile enhances the details of each tile. The model can identify various areas in the image and enhance the details of each area, adding additional detail where necessary. The detail enhancement process can be expressed as:
@@
Step 2: RealESRGAN Super-Resolution Reconstruction Based on ControlNet-Tile, we further employed RealESRGAN for higher precision super-resolution reconstruction. RealESRGAN is an advanced super-resolution method based on generative adversarial networks, capable of enhancing resolution while preserving natural details and visual consistency. The super-resolution reconstruction process can be expressed as:
@@
Step 3: ControlNet-Tile Re-Optimization After RealESRGAN processing, we applied ControlNet-Tile once more for additional refinement and optimization. The purpose of this second application of ControlNet-Tile is to further enhance the details and quality of the image at the increased resolution. The final optimized image $I_{f}inal$ is represented as:
@@
Multi-View Image Acquisition. After obtaining the high-resolution panorama, we need to extract images from specific viewpoints for point cloud reconstruction. We select a suitable set of camera poses $P_{i}\in\mathbb{R}^{4\times 4}$ based on the requirements of point cloud reconstruction. The selection of these camera poses is based on geometric and visual coverage principles, ensuring that the chosen perspectives comprehensively cover important details and areas of the panorama. Specifically, we use an optimization-based viewpoint selection algorithm that considers the distribution of viewpoints, parallax effects, and coverage of reconstruction areas. The camera pose $P_{i}$ is represented as:
@@
By applying the above methods, we can generate high-quality 360-degree panoramas from input text prompts, providing high-resolution and detailed input data for subsequent point cloud reconstruction. Through precise perspective transformations, we obtained a set of viewpoint images that satisfy the requirements of point cloud reconstruction. These viewpoint images serve as a solid foundation for the subsequent point cloud reconstruction process.
Point Cloud Construction with Gaussian Splatting
@@
Point Cloud Initialization. The initial point cloud must be generated from multi-view images to enable reconstruction using 3D Gaussian Splatting (Kerbl et al. 2023). To generate a 360-degree spatial point cloud from multi-view images, a rendering trajectory centered around the observation viewpoint is required. This trajectory should cover the bottom, middle, and top of the spatial area. Using the LucidDreamer (Chung et al. 2023) framework, we incorporate additional camera viewpoints to create a more accurate spatial point cloud. For the set of multi-view images $\mathbf{I_{1}},\mathbf{I_{2}},...,\mathbf{I_{n}}$ derived from the panoramic image, each image $\mathbf{I_{i}}$ has a corresponding position $\mathbf{P_{i}}$. Using the monocular depth estimation model ZoeDepth (Bhat et al. 2023), we measure the depth map $\mathbf{D_{i}}$ for each $\mathbf{I_{i}}$, forming the depth map set $\mathbf{D_{1}},\mathbf{D_{2}},...,\mathbf{D_{n}}$. We initialize the point cloud using a cyclic generation approach. The camera’s intrinsic matrix and the extrinsic matrix of $\mathbf{I_{0}}$ are denoted as $\mathbf{K}$ and $\mathbf{P_{0}}$, respectively. The initial point cloud for a single viewpoint is derived using the following formula:
@@
Rendering with Gaussian Splatting. 3D Gaussian Splatting is a technique that can quickly and efficiently model 3D scenes. After generating the point cloud, we use it to train a 3D Gaussian Splatting model. The generated point cloud $\Omega$ serves as the initial Structure from Motion (SfM) points. Initializing with $\Omega$ accelerates network convergence and encourages the network to focus on generating detailed representations. We utilize multi-view image slices as training images. Additionally, we apply the image view augmentation technique from LucidDreamer. Since the input multi-view images may not cover the complete 3D space, we project the $\Omega$ point cloud into 2D space from different camera viewpoints to increase the number of supervisory images. The formal representation are as follows:
@@
Figure 3: The visualization of SceneDreamer360. The Panoramic Image on the left represents the high-resolution enhanced panorama generated from a specific text input. The Multi-view Image shows different perspective images derived from the panorama, while the Output on the right displays images rendered from the generated point cloud at new viewpoints. As shown in the figure, SceneDreamer360 is capable of producing high-quality, detailed, and spatially consistent 3D scenes.
Experiment
@@
Implementation Details. Our method is implemented using PyTorch. In the first stage, we generate panoramic images using a modified PanFusion model. We fine-tune this model on the Habitat Matterport Dataset, incorporating MLP and LoRA layers to produce realistic 3D panoramic images. A three-step super-resolution reconstruction process, including ControlNet-Tile and RealESRGAN, enhances the resolution of these images to $3072\times 6144$. In the second stage, we adapt and modify the point cloud reconstruction component from LucidDreamer. For evaluation, we use GPT-4 to generate a set of prompts that cover a wide range of scene types. We assess the quality of our outputs using metrics such as CLIP-Score, PSNR, and others. Our experiments are conducted on an NVIDIA A100-80G GPU.
Baselines. We compare our SceneDreamer360 with the following baselines.
  • TextToRoom (Höllein et al. 2023)employs a mesh-based 3D representation, directly extracting meshes from inpainted RGBD images to represent watertight indoor environments.
  • LucidDreamer (Chung et al. 2023) projects outpainted RGBD images onto a point cloud and then supervises the optimization of 3DGS by projecting multiple images from this point cloud. However, since LucidDreamer cannot directly generate 3D scenes from prompts, we utilize a diffusion model to generate condition images, allowing LucidDreamer to create scenes from text input.
Evaluation Metrics. Since there are no established evaluation metrics for 3D scene point clouds, we assess our method using metrics from prior work. We calculate the CLIP-Score (Hessel et al. 2021) to measure text-image consistency, employ CLIP-IQA (Wang, Chan, and Loy 2023) to evaluate image quality from a single viewpoint, and use PSNR, SSIM, and LPIPS to assess the quality of the rendered images.
Figure 4: The visualization comparison of the proposed SceneDreamer360 between current methods. The images on the left show the new viewpoint renderings, while the images on the right display the spatial views of the generated 3D scenes.
Qualitative Result
The point cloud images generated by our method using different prompts are shown in Fig. 3. The method successfully generates 3D spatial scenes consistent with the text prompts. Additionally, part of the output demonstrates that images can be rendered from arbitrary trajectory views, showcasing the generalizability of our method. We visually compared our method with LucidDreamer and Text2Room using identical spatial text prompts, as shown in Fig. 4. Our method outperforms the others by producing more consistent, complete, and detailed 3D spatial scene point clouds. The use of 3DGS enables superior detail and continuous object surfaces, surpassing Text2Room in clarity. Additionally, our Stage 1 panoramic optimization captures more spatial details, and our designed rendering trajectory generates more comprehensive scenes than LucidDreamer. For instance, LucidDreamer’s point clouds are incomplete with visible voids, while Text2Room struggles with clarity, particularly in floor details. In contrast, SceneDreamer360 delivers consistent and aesthetically pleasing indoor scenes that accurately reflect the prompt, demonstrate the superiority of our approach.
@@
Text2Room(Höllein et al. 2023) - - - - 19.29 0.728 0.206
LucidDreamer(Chung et al. 2023) 0.2110 0.9437 0.3230 0.4757 29.56 0.922 0.039
@@
Table 1: Comparison of image and render quality across different methods.
Results Comparision
To objectively evaluate the generation quality, we utilized metrics such as CLIP-Score, CLIP-IQA, PSNR, SSIM, and LPIPS. Table 1 presents the quantitative evaluation results. For generated image quality, SceneDreamer360 outperformed LucidDreamer in terms of CLIP-Score, CLIP-IQA-sharp, CLIP-IQA-colorful, and CLIP-IQA-resolution, achieving better scores compared to LucidDreamer’s scores. Due to the significant variations in perspective and incomplete scenes generated by Text2Room, we did not measure its image quality. However, as illustrated in Fig. 4, the consistency of the scenes generated by Text2Room is notably poor, often resulting in unrealistic scenarios. These results demonstrate the significant advantage of our method in generating high-quality and consistent 3D multiview images. For render quality of the 360-degree scenes, we evaluated the generated images by calculating the average of metrics. Our method outperformed others in terms of PSNR, SSIM, and LPIPS, demonstrating the effectiveness of our approach.
Figure 5: The ablation study of SceneDreamer360.
Ablation Study
We conducted ablation experiments on two critical modules: panoramic image enhancement and point cloud fusion, to investigate the effectiveness of our method. The key ablation configurations included: “Standard”, “W/o enhance”, “W/o alignment”, and “W/o enhance & alignment”. As shown in Fig. 5, removing the enhancement module resulted in a loss of finer details, while omitting the alignment module led to noticeable visual discontinuities. The reconstruction without both modules produced worst results, losing both detail and spatial consistency. Additionally, we measured the quantitative performance of the reconstructed images across different configurations, as presented in Table 2. This indicates that removing the alignment module causes a more significant decline in metrics than removing the enhancement module, highlighting the alignment module’s crucial role in maintaining consistent spatial generation.
@@
Table 2: Quantitative comparison of different settings.
Conclusion

### v2

@@ III-C1 ControlNet-Tile Super-Resolution Reconstruction
The ControlNet-Tile model divides the panorama into distinct tiles through semantic recognition. Each tile is analyzed by the model to determine which details to add and which parts to repair. Based on tiling, ControlNet-Tile enhances the details of each tile. The model can identify various areas in the image and enhance the details of each area, adding additional detail where necessary. The detail enhancement process can be expressed as:
@@
III-C2 RealESRGAN Super-Resolution Reconstruction
Based on ControlNet-Tile, we further employed RealESRGAN for higher precision super-resolution reconstruction. RealESRGAN is an advanced super-resolution method based on generative adversarial networks, capable of enhancing resolution while preserving natural details and visual consistency. The super-resolution reconstruction process can be expressed as:
@@ III-C2 RealESRGAN Super-Resolution Reconstruction
III-C3 ControlNet-Tile Re-Optimization
After RealESRGAN processing, we applied ControlNet-Tile once more for additional refinement and optimization. The purpose of this second application of ControlNet-Tile is to further enhance the details and quality of the image at the increased resolution. The final optimized image $I_{f}inal$ is represented as:
@@ III-C3 ControlNet-Tile Re-Optimization
III-D Multi-View Image Acquisition
After obtaining the high-resolution panorama, we need to extract images from specific viewpoints for point cloud reconstruction. We select a suitable set of camera poses $P_{i}\in\mathbb{R}^{4\times 4}$ based on the requirements of point cloud reconstruction. The selection of these camera poses is based on geometric and visual coverage principles, ensuring that the chosen perspectives comprehensively cover important details and areas of the panorama. Specifically, we use an optimization-based viewpoint selection algorithm that considers the distribution of viewpoints, parallax effects, and coverage of reconstruction areas. The camera pose $P_{i}$ is represented as:
@@ III-D Multi-View Image Acquisition
Based on the camera poses, the e2p algorithm generates perspective images from the panorama. These perspective images retain the high resolution and details of the panorama while introducing parallax between views, providing rich perspective information for subsequent point cloud reconstruction.
By employing these methods, we can effectively generate high-quality, 360-degree panoramas directly from text prompts. This approach provides high-resolution and detailed input data, forming an optimal basis for the subsequent point cloud reconstruction. Using precise perspective transformations, we produce a series of viewpoint images that are well-suited to the requirements of point cloud reconstruction. These images not only capture various angles but also retain the visual fidelity and spatial coherence needed for constructing accurate 3D scenes. Consequently, they serve as a solid foundation for the point cloud reconstruction process, enhancing the realism and consistency of the final generated scene.
III-E Point Cloud Construction with Gaussian Splatting
@@
III-E1 Point Cloud Initialization
The initial point cloud must be generated from multi-view images to enable reconstruction using 3D Gaussian Splatting [8]. To generate a 360-degree spatial point cloud from multi-view images, a rendering trajectory centered around the observation viewpoint is required. This trajectory should cover the bottom, middle, and top of the spatial area. Using the LucidDreamer [6] framework, we incorporate additional camera viewpoints to create a more accurate spatial point cloud. For the set of multi-view images $\mathbf{I_{1}},\mathbf{I_{2}},...,\mathbf{I_{n}}$ derived from the panoramic image, each image $\mathbf{I_{i}}$ has a corresponding position $\mathbf{P_{i}}$. Using the monocular depth estimation model ZoeDepth [62], we measure the depth map $\mathbf{D_{i}}$ for each $\mathbf{I_{i}}$, forming the depth map set $\mathbf{D_{1}},\mathbf{D_{2}},...,\mathbf{D_{n}}$. We initialize the point cloud using a cyclic generation approach. The camera’s intrinsic matrix and the extrinsic matrix of $\mathbf{I_{0}}$ are denoted as $\mathbf{K}$ and $\mathbf{P_{0}}$, respectively. The initial point cloud for a single viewpoint is derived using the following formula:
@@ III-E1 Point Cloud Initialization
III-E2 Rendering with Gaussian Splatting
3D Gaussian Splatting is a technique that efficiently models 3D scenes by leveraging a fast and accurate point-based representation. Following the generation of the initial point cloud, we use this point cloud to train a 3D Gaussian Splatting model. The generated point cloud, denoted as $\Omega$ , serves as the initial Structure from Motion (SfM) points, which accelerates network convergence and encourages the model to focus on producing detailed and realistic representations of the scene. To train the model, we utilize multi-view image slices as training images and incorporate the image view augmentation technique from LucidDreamer to enrich the data. Since the input multi-view images may not fully cover the entire 3D space, we project the $\Omega$ point cloud into 2D from various camera viewpoints, effectively increasing the number of supervisory images and expanding spatial coverage. This augmentation process helps to provide more comprehensive training data, ensuring better spatial consistency and detail across the 3D scene.
@@ III-E2 Rendering with Gaussian Splatting
Figure 3: The visualization comparison of the proposed SceneDreamer360 between current methods. The images on the left display the new viewpoint renderings, demonstrating how each method handles various perspectives. In contrast, the images on the right present the spatial views of the generated 3D scenes, allowing a direct comparison of the spatial consistency and overall structure of the reconstructions.
IV Experiments
@@
IV-A Implementation Details
In the first stage, we generate high-quality panoramic images using a modified PanFusion model, which has been fine-tuned on the Habitat Matterport Dataset. To enhance its capacity for producing realistic 3D panoramic images, we incorporate MLP and LoRA layers into the model. This fine-tuning enables the generation of more visually coherent and lifelike panoramas. We further apply a three-step super-resolution process, using ControlNet-Tile and RealESRGAN, which enhances the resolution of these panoramas to $3072\times 6144$, preserving fine details and enriching the image quality for subsequent stages.
In the second stage, we adapt and modify the point cloud reconstruction component from LucidDreamer. This adaptation allows us to effectively convert the high-resolution panoramic images into detailed 3D point clouds, ensuring consistency with the input scene and text prompt. This two-stage process of panorama generation and point cloud reconstruction is designed to maximize visual fidelity and structural accuracy in the final 3D scenes.
Our method is implemented using PyTorch. We obtain camera viewpoints from given camera poses and use the e2p algorithm to extract corresponding multi-view images from the panorama. The specific implementation method is as follows:
[…]
test.py112
For evaluation, we generate a diverse set of prompts using GPT-4, covering a broad spectrum of scene types to comprehensively test our model’s versatility. We then assess the quality of our outputs using multiple metrics, including CLIP-Score, PSNR, and others, to evaluate both visual similarity and fidelity to the input text prompts. All experiments are conducted on an NVIDIA A100-80G GPU, ensuring sufficient computational power for handling high-resolution image processing and point cloud generation.
Figure 4: The Panoramic Image on the left represents the high-resolution enhanced panorama generated from a specific text input, showcasing the detailed and cohesive scene captured in 360 degrees. The Multi-view Image in the center displays various perspective images derived from this panorama, illustrating the scene’s adaptability across different viewpoints. Finally, the Output on the right presents images rendered from the generated point cloud at new viewpoints, highlighting SceneDreamer360’s ability to maintain high quality and spatial consistency in the 3D reconstructions. As shown in the figure, SceneDreamer360 effectively produces detailed, visually coherent, and spatially consistent 3D scenes.
TABLE I: Comparison of image and render quality across different methods.
@@ IV-A Implementation Details
Text2Room [5] - - - - 19.29 0.728 0.206
LucidDreamer [6] 0.2110 0.9437 0.3230 0.4757 29.56 0.922 0.039
@@
TABLE II: Quantitative comparison of different settings.
@@
IV-B Baselines
We compare our SceneDreamer360 with the following state-of-the-art baselines:
  • TextToRoom [5] employs a mesh-based 3D representation, directly extracting meshes from inpainted RGBD images to represent watertight indoor environments.
  • LucidDreamer [6] projects outpainted RGBD images onto a point cloud and then supervises the optimization of 3DGS by projecting multiple images from this point cloud. However, since LucidDreamer cannot directly generate 3D scenes from prompts, we utilize a diffusion model to generate condition images, allowing LucidDreamer to create scenes from text input.
Evaluation Metrics As there are currently no established evaluation metrics specifically for 3D scene point clouds, we assess our method using metrics adapted from prior work in related areas. We calculate the CLIP-Score [63] to measure text-image consistency, allowing us to evaluate how well the generated images align with the input text prompts. To assess the visual quality of single viewpoint renderings, we use CLIP-IQA [64], which provides an image quality evaluation aligned with human perception. Additionally, we employ PSNR, SSIM, and LPIPS to evaluate the fidelity and perceptual quality of the rendered images. These metrics collectively enable a well-rounded assessment, capturing both the accuracy of text-image correspondence and the visual quality of the generated 3D scenes.
IV-C Qualitative Result
The point cloud images generated by our method using various prompts are displayed in Fig. 4. These results demonstrate that our method effectively creates 3D spatial scenes that are closely aligned with the input text prompts. Additionally, we show that the generated images can be rendered from arbitrary viewpoint trajectories, highlighting the generalizability and flexibility of our approach. To further evaluate our method, we visually compared the output with LucidDreamer and Text2Room using identical spatial text prompts, as illustrated in Fig. 3. Our method surpasses these alternatives by producing more consistent, complete, and detailed 3D spatial scene point clouds. Specifically, the integration of 3D Gaussian Splatting (3DGS) enables our model to capture finer details and maintain continuous object surfaces, resulting in sharper, more defined images compared to Text2Room, which struggles with clarity and fine details.
Our Stage 1 panoramic optimization also contributes significantly by capturing extensive spatial details, while our tailored rendering trajectory produces comprehensive, cohesive scenes. In contrast, LucidDreamer frequently generates point clouds that are incomplete, with visible voids that compromise the spatial coherence of the scene. Text2Room, on the other hand, often fails to achieve clarity, particularly in regions such as floors and other large surfaces, leading to a lack of detail. In comparison, SceneDreamer360 consistently delivers aesthetically pleasing and well-detailed indoor scenes that accurately reflect the text prompts. The results underscore the superiority of our approach in achieving high-quality, text-consistent, and spatially complete 3D scene reconstructions.
IV-D Results Comparision
To objectively assess the quality of the generated images, we employ a range of evaluation metrics, including CLIP-Score, CLIP-IQA, PSNR, SSIM, and LPIPS. Table I presents the quantitative results of our evaluation. In terms of generated image quality, SceneDreamer360 consistently outperforms LucidDreamer across multiple metrics, achieving higher scores in CLIP-Score, CLIP-IQA-sharp, CLIP-IQA-colorful, and CLIP-IQA-resolution. These results indicate that our method produces sharper, more colorful, and higher-resolution images that reflect the text prompts with greater fidelity compared to LucidDreamer. Due to the significant perspective variations and incomplete scenes produced by Text2Room, we do not include it in our quantitative image quality measurements. As shown in Fig. 3, however, the scenes generated by Text2Room display poor consistency and often result in unrealistic or fragmented visual scenarios. This highlights the clear advantage of SceneDreamer360 in generating coherent and realistic multiview 3D scenes.
To evaluate the overall render quality of the 360-degree scenes, we calculate the average values for PSNR, SSIM, and LPIPS across the generated images. SceneDreamer360 consistently outperforms other methods, achieving higher PSNR and SSIM scores, which indicate superior image fidelity and structural similarity. Additionally, our method achieves lower LPIPS scores, reflecting enhanced perceptual quality in the generated images. These findings underscore the effectiveness of our approach in producing high-quality, consistent, and visually appealing 3D multiview images, thereby setting a new benchmark for text-driven 3D scene generation.
Figure 5: The ablation study of SceneDreamer360.
IV-E Ablation Study
We conduct ablation experiments on two critical modules—panoramic image enhancement and point cloud alignment—to evaluate their individual contributions to the effectiveness of our method. The primary configurations tested are: “Standard” (using both modules), “W/o Enhance” (without the enhancement module), “W/o Alignment” (without the alignment module), and “W/o Enhance & Alignment” (without both modules).
As illustrated in Fig. 5, the absence of the enhancement module results in a noticeable loss of fine details, indicating that this module is essential for preserving high-resolution features. When the alignment module is removed, we observe significant visual discontinuities, underscoring its role in maintaining spatial coherence across the generated scenes. Notably, the configuration without both modules produces the poorest results, lacking both detail and spatial consistency, which detracts from the realism of the reconstructed 3D scenes.
To further quantify these effects, we measure the performance of the reconstructed images across the different configurations, as shown in Table II. The results indicate that removing the alignment module leads to a more substantial decline in metrics compared to the removal of the enhancement module, highlighting the alignment module’s critical role in ensuring consistent and coherent spatial generation. These findings demonstrate that both modules are integral to our method, with alignment playing a particularly crucial role in achieving high-quality 3D scene reconstruction.
V Conclusion